Journalist Report – October 28th

Monday 28th October Sol 1

After a freezing night (literally) we awake to a cold but clear day. The mountain to the south (Henry Mountains) is wreathed in cloud, and shows a glazing of snow. Today is our first day ‘on Mars’! This takes a bit of psychological adjustment.

We congratulate each other on a successful ‘journey’, and get on with the day. This is mainly about bedding in systems: getting the radios working, making equipment for the micrometeorite experiment, setting up the fire blanket for the kitchen.

The radios were a bit of a mess, with a number not working in different ways, but we were able to get a working minimum set functioning eventually.

Today we are expecting our 5th team member, David, to arrive, or ‘descend in the latest shuttle’. David is very experienced in the MDRS, so will be able to train us up in the use of the suits, etc.

The afternoon shift was spent working on equipment for the micrometeorite expedition, tying and marking string in 1 metre lengths for the grid to be set up later in the field.
After another delicious re-constituted meal, we move onto the CapCom reports and wind up the day.

Operations Report – October 28th

Name of person filing report: David Mateus
Non-nominal systems: Opportunity
Notes on non-nominal systems: Its brakes do not work
Generator: run
Hours run: 8.5
From what time last night: 18:00
To what time this morning: 2:30
List any additional daytime hours when the generator was run: none
Solar— SOC % (Before generator is run at night): 84%
Diesel Reading –75 %
Station Propane Reading – 88 %
Ethanol Free Gasoline – N/A
Water (loft tank): 47 gallons
Water Meter: 1454197 units
Water (static tank) – 350 gallons
Static to Loft Pump used – Yes
Water in GreenHab – 0 gallons
Water in ScienceDome: 0 gallons
Toilet tank emptied: yes

Deimos rover used: It is not in the hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: No
Hours: (before EVA): 105.5 h
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging):100%
Currently charging: Yes

Opportunity rover used: No
Hours: 65.5 h
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 111.5
Beginning charge: 100%
Ending charge:100%
Currently charging: Yes

Notes on rovers: No

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) No
Reason for use: None
Oil Added? No
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? Used by assistant directors to go to town

CrewCar used and why, where? Used to go to Grand Junction

General notes and comments: None
Summary of internet: functional
Summary of suits and radios: No used
Summary of Hab operations: All systems are nominal
Summary of GreenHab operations: No used
Summary of ScienceDome operations: All systems are nominal, used for experiment preparations
Summary of RAM operations: No used
Summary of any observatory issues: Re-centered the focuser by the assistant directors
Summary of health and safety issues: All the crew are healthy
Questions, concerns and requests to Mission Support: We need new batteries for six radios.

Sol Summary – October 28th


Summary Title: Touchdown on Mars

Author’s name: Guy Murphy

Mission Status: The crew woke up on Mars this morning, this being our first day in full simulation mode.

Sol Activity Summary: The morning was dedicated to preparing and commencing use of the radios, as well as other minor maintenance, cleaning and administrative tasks. In the afternoon our 5th crew member David Mateus arrived. Crew geologist Andrew Wheeler continued preparation of equipment for the micro-meteorite collection study. Three cooked meals were prepared today.

Look Ahead Plan: Our goal is to commence our EVA schedule tomorrow, with David Mateus leading us on our first EVAs to learn about the procedures and equipment.

Anomalies in work: Not all the radios were found to be working.

Weather: Very cold overnight, clear and sunny daytime sky.

Crew Physical Status: All crew in good health.

EVA: Not commenced today.

Reports to be filed: Sol Summary, Operations Report, EVA Requests (2), Journalists Report, Science Report, Days Photos, Crew Car Collection Report, Mission Patch.

Support Requested: New batteries were requested to supplement the existing limited supply of working batteries for future crews.

Copyright © The Mars Society. All rights reserved. | Main Site